OpenStudy (anonymous):

Find 3rd term of the expansion of (2 + 3x)^8

OpenStudy (anonymous):

is there a way of doing this other than writing the expansion term by term?

OpenStudy (anonymous):

Is your expansion around x=0?

OpenStudy (anonymous):

If \[ f(x)=(2 + 3x)^8\\ f'(x) =24 (3 x+2)^7\\ f''(x)=504 (3 x+2)^6\\ \frac {f''(0)}{2!}=16128\\ \] The third term is: \[16128 x^2 \] If you expand f(x), you get \[ f(x) = 6561 x^8+34992 x^7+81648 x^6\\+108864 x^5+90720 x^4+48384 x^3+16128 x^2\\+3072 x+256 \]

OpenStudy (amistre64):

you seem to have found the anti-3rd term :)

OpenStudy (amistre64):

i spose that depends on how you define a proper polynomial eh ...

OpenStudy (anonymous):

the answer given in the book was 16128x^2 - i suppose it depends which way you order the terms thanx guys

OpenStudy (anonymous):

The third term in a power series is of a functiom f at zero is \[ \frac {f''(0)}{2!} x^2 \] Since the series starts as \[ \frac {f(0)}{0!} x^0+ \frac {f'(0)}{1!} x^1 + \frac {f''(0)}{2!} x^2+ \cdots \]